Can you show me how to calculate the Profit in this one? Help me please!
3ムS 29 34 45 455 55 ad
Answer #1

Ans: Before solving this question, one needs to understand how can we calculate profit.

Profit = Total Revenue - Total Cost

From the given ques, we can estimate that total cost will be equal to the wages given to the laborers, i.e., TC= L*W. Similarly. total revenue can be estimated by the price of the output produced, i.e, TR= P*Q.

Using the above formulae, given wages=$40 and price= $10

Q L TC (L*W) TR (P*Q) PROFIT (TR-TC)
10 1 40 100 60
19 2 80 190 110
27 3 120 270 150
34 4 160 340 180
40 5 200 400 200
45 6 240 450 210
49 7 280 490 210
52 8 320 520 200
54 9 360 540 180
55 10 400 550 150
385 55 2200 3850 1650

Total Profit= Total Revenue- Total Cost

or, 1650= 3850-2200